OpenStudy (anonymous):

find all real solutions to 2sin^2x+3sinx+1

OpenStudy (phi):

I assume you mean \[ 2\sin^2x+3\sin x+1=0 \] ? If so, temporarily rename sin(x) as y so that you have the equation \[ 2y^2 +3y+1=0 \] can you factor this to find the y values?

OpenStudy (anonymous):

Yeah sorry! i meant that

OpenStudy (phi):

can you factor the quadratic 2y^2 +3y+1=0 ?

OpenStudy (anonymous):

yes! sorry i got lost